# Parameter-Efficient Fine-Tuning of Large Language Models for Unit Test Generation: An Empirical Study

### v1

@@ V Experimental Results
V-A RQ1: Tuning method performance
We present the performance results of the unit test generation for the various models and tuning methods in Table IV.
TABLE IV: Comparison of syntactical validity and CodeBLEU scores from experiments using different tuning methods across various models on testing split of Methods2Testsmall and HumanEval-Xjava datasets.
Model Method Trainable
@@ V-A RQ1: Tuning method performance
Methods2Testsmall HumanEval-Xjava
Valid syntax
CodeBLEU
@@
Instr Cov
Branch Cov
CodeGen-350M-multi None 0 95.43% 0.2170 100% 0.3608 0.0671 97.33% 89.77%
Full fine-tuning 304.23M 97.87% 0.2988 100% ( 0.3293 ) ( 0.0366 ) 100% ( 83.33% )
LoRA 1.31M ( 95.22% ) 0.2553 100% 0.3907 0.0671 98.69% ( 89.29% )
(IA)3 0.14M 95.53% 0.2266 100% ( 0.3583 ) ( 0.0549 ) 97.69% 94.32%
Prompt tuning 0.02M 96.03% 0.2208 100% ( 0.3290 ) ( 0.0427 ) ( 96.56% ) 91.67%
CodeGen2-1B None 0 0% 0 0% 0 0 0% 0%
Full fine-tuning 1,015.31M 76.73% 0.1474 5.49% 0.0359 0 0% 0%
LoRA 2.10M 41.16% 0.0484 8.54% 0.0117 0 0% 0%
(IA)3 0.23M 1.52% 0.2553 0% 0 0 0% 0%
Prompt tuning 0.04M 66.63% 0.2568 7.93% 0.2547 0 0% 0%
StarCoder2-3B None 0 85.23% 0.1543 100% 0.4264 0.3152 9.89% 85.67%
Full fine-tuning 3,030.37M 96.71% 0.2786 100% 0.4969 0.3494 99.40% 86.37%
LoRA 4.55M 97.11% 0.2901 100% ( 0.4169 ) 0.3675 99.19% 58.84%
(IA)3 0.47M 87.43% 0.2513 100% ( 0.4250 ) ( 0.2744 ) 99.67% ( 83.81% )
Prompt tuning 0.06M 86.43% 0.1742 100% 0.4309 ( 0.2470 ) 99.6% ( 75.85% )
CodeGen2-3.7B None 0 0% 0 0% 0 0 0% 0%
Full fine-tuning 3,641.17M 50.51% 0.1006 73.78% 0.2621 0 0% 0%
LoRA 4.19M 52.24% 0.0997 40.24% 0.1384 0 0% 0%
(IA)3 0.46M 0% 0 0% 0 0 0% 0%
Prompt tuning 0.08M 23.50% 0.2562 0% 0 0 0% 0%
CodeLlama-7B None 0 97.66% 0.3107 99.39% 0.4861 0.3293 98.33% 84.46%
Full fine-tuning 6,607.41M ( 96.44% ) ( 0.3012 ) 100% 0.4994 0.3373 98.95% 86.37%
LoRA 8.39M ( 97.36% ) 0.3277 99.39% ( 0.4291 ) ( 0.3129 ) 99.61% ( 72.47% )
(IA)3 0.61M ( 97.05% ) ( 0.3011 ) 100% ( 0.4802 ) ( 0.3232 ) 98.77% 84.72%
Prompt tuning 0.08M ( 95.93% ) ( 0.2885 ) 99.39% ( 0.4617 ) ( 0.2761 ) 98.38% ( 82.25% )
CodeGen2-7B None 0 96.95% 0.2848 100% 0.4736 0.2256 98.31% 81.45%
Full fine-tuning 6,862.87M 97.56% 0.3107 100% ( 0.4398 ) ( 0.1280 ) 99.75% ( 70.00% )
LoRA 8.39M 97.87% 0.3164 100% ( 0.4636 ) ( 0.2073 ) ( 98.06% ) ( 75.35% )
(IA)3 0.92M 97.36% 0.2904 100% 0.4898 ( 0.1829 ) 98.55% ( 79.50% )
Prompt tuning 0.08M ( 96.64% ) ( 0.2775 ) 100% ( 0.4407 ) ( 0.2012 ) 99.10% ( 69.40% )
StarCoder2-7B None 0 84.13% 0.1610 100% 0.4027 0.3758 99.07% 83.16%
Full fine-tuning 7,173.92M 97.21% 0.3009 100% 0.4389 ( 0.3675 ) 99.15% 90.80%
LoRA 7.34M 96.91% 0.3068 100% 0.5179 ( 0.3394 ) 99.35% 87.06%
(IA)3 0.75M 94.83% 0.2903 100% 0.4213 ( 0.3697 ) 99.40% 88.39%
Prompt tuning 0.09M ( 83.03% ) 0.3030 100% 0.5057 ( 0.3476 ) 99.38% 86.02%
StarCoderBase None 0 84.63% 0.1563 98.78% 0.4338 0.2963 99.07% 81.48%
Full fine-tuning 15,517.46M 96.81% 0.3123 100% 0.4830 0.3293 99.16% ( 75.20% )
LoRA 8.03M 95.71% 0.3152 98.78% ( 0.3905 ) 0.2963 99.07% ( 73.89% )
(IA)3 1.24M 84.63% ( 0.1553 ) 98.78% 0.4344 ( 0.1562 ) ( 98.72% ) 81.48%
Prompt tuning 0.12M 85.73% ( 0.1518 ) ( 78.05% ) ( 0.2315 ) 0.3025 99.76% ( 67.62% )
StarCoder2-15B None 0 85.43% 0.1898 100% 0.3724 0.4085 98.93% 87.69%
Full fine-tuning 15,655.90M 97.90% 0.3323 ( 99.39% ) 0.4886 0.3758 99.52% 81.3%
LoRA 12.12M 97.01% 0.3272 100% 0.4633 0.4146 98.95% 82.88%
(IA)3 1.25M 85.43% 0.1901 100% 0.3725 0.4578 99.57% 87.89%
Prompt tuning 0.12M 97.60% 0.3133 100% 0.5352 0.3939 99.32 % 82.89 %
CodeGen2-16B None 0 97.87% 0.2784 100% 0.4779 0.2012 98.66% 80.56%
Full fine-tuning 16,032.16M ( 97.56% ) 0.3383 ( 98.17% ) ( 0.3774 ) ( 0.1180 ) 99.52% ( 78.07% )
LoRA 13.37M 98.68% 0.3186 100% ( 0.4714 ) 0.2012 98.66% 82.06%
(IA)3 1.46M 97.87% 0.2790 100% 0.4780 0.2134 ( 97.46% ) 80.56%
Prompt tuning 0.08M 97.97% 0.2954 100% ( 0.4679 ) 0.2195 ( 98.62% ) ( 71.07% )
  • Bold: best-performing training method per model. ( Parentheses ): decreased performance compared to baseline. Red: $<$ 50% syntactical valid samples. Underline: Other notable results (see in Section V-A1).
V-A1 Syntactic Correctness
@@
As shown in Table IV, the percentage of syntactically correctly generated unit tests is relatively high across the board, with $>$ 80% for all except the smaller CodeGen2-1B and CodeGen2-3.7B models. These two cannot produce any valid unit test with the base model. However, after tuning, both models are able to produce up to 76.73% and 52.24% syntactically valid tests, respectively. The notable findings (marked with underlines in Table IV) are as follows: 1) The increase of 66.63% using prompt tuning of CodeGen2-1B, tuning only 0.02‰ (0.04 million out of 1 billion parameters) of the parameters. 2) The increase of 52.24% using LoRA on CodeGen2-3.7B, tuning only 0.1% (4.19 million out of 3.7 billion parameters).
@@
Further, data in Table IV illustrate that: 1) Full fine-tuning achieves higher percentages of syntactically correct unit tests than the baseline. The median percentage of syntactically valid tests for the full fine-tuning is 97.01%, which is 13.69% higher than the value of the baseline (85.33%). 2) There is a weak tendency that the more trainable parameters, the more samples are generated syntactically correctly. The median percentage of syntactically valid tests for the full fine-tuning is 97.01%, which is slightly higher than LoRA (96.96%), (IA)3 (91.13%), and prompt tuning (91.18%). However, considering that full fine-tuning often tunes parameters several thousands more times than PEFT methods, it is probably not wise to use full fine-tuning to improve the syntactic correctness of the unit tests if cost-effectiveness is a priority.
V-A2 Similarity
The Methods2Testsmall dataset CodeBLEU column in Table IV shows the CodeBLEU scores of the baseline, full fine-tuning, and PEFT methods. The notable results are: 1) LoRA archives the highest CodeBLEU scores in five of the ten models across different sizes. LoRA is also the only method that can positively increase the CodeBLEU score of the CodeLlama-7B model. 2) Full fine-tuning is nearly as performant as LoRA. For three of the ten models, full fine-tuning scores the best. 3) (IA)3 performs generally poorly compared to others, often resulting in very low CodeBLEU scores. 4) Prompt tuning shows varying results, from being the worst performing on CodeLlama-7B to being the best performing on CodeGen2-1B.
V-A3 Coverage
@@
From the pass@1 column in Table IV, the average mean pass@1 rate is 0.2332. Of the tests that pass, the code coverage is generally high, as can be seen from the two last columns in Table IV. The exception is where the model is not able to generate much syntactically valid code. This results in some coverage results being zero, e.g. CodeGen2-1B and CodeGen2-3.7B. With an average median instruction coverage of 98.81% and branch coverage of 78.55%, this shows that if the models are able to create a runnable test, they are likely to cover a significant portion of the code. This high level of coverage indicates that the generated tests are effective at exercising the codebase, potentially identifying edge cases, and ensuring robustness. The results also show that LoRA and full fine-tuning outperform other methods in several cases. Although (IA)3 and prompt tuning perform best in some cases, the differences between their performance and others are insignificant.
V-A4 Resilience against catastrophic forgetting
@@
The results in the last five columns of Table IV show the evaluation of various PEFT tuning methods on the Java subset of HumanEval-X after they have been trained on the Methods2Testsmall dataset. As expected, we can see from the results that there are some cases of degradation (marked using parentheses in Table IV) when comparing the results to the baseline. This applies to both full fine-tuning and PEFT. However, the levels of degradation are, for the most part, trivial. We also see several cases, e.g., StarCoder2-15B, where PEFT tuning actually improves the CodeBLEU scores. This shows that neither tuning method is very susceptible to catastrophic forgetting, and PEFT methods might pose similarly good transfer learning capabilities to that of full-fine-tuning. Summary for RQ1: Full fine-tuning and PEFT methods could improve the syntactic correctness of the generated unit tests. PEFT methods could be better choices than full fine-tuning if cost-effectiveness to improve the syntactic correctness of the generated unit tests is a concern. Full fine-tuning and LoRA could significantly boost the quality of the generated unit tests measured using CodeBLU, with LoRA even surpassing full fine-tuning in several cases. Further, PEFT methods are mostly resilient against catastrophic forgetting.
V-B RQ2: PEFT training resource utilization effectiveness
$10^{6}$$10^{8}$$10^{10}$rainable parametersCodeGen-350M-multiCodeGen2-1BStarCoder2-3BCodeGen2-3.7BCodeLlama-7BCodeGen2-7BStarCoder2-7BStarCoderBaseStarCoder2-15BCodeGen2-16BFull fine-tuningLoRA(IA)3Prompt tuning
Figure 2: Comparison of the number of trainable parameters for each tuning method. Models are sorted by size.

### v2

@@ 5 Experimental Results
5.1 RQ1: Tuning method performance
We begin by describing the performance results for each evaluation metric on the Methods2Testrunnable dataset, as shown in Section 5.1(b). We then explain the results to test the catastrophic forgetting effects on code generation (shown in Section 5.1(c)) using the HumanEval-X dataset.
Evaluation metrics experiment results using different tuning methods across various models.
(a) (b) (c)
Method Trainable
@@ 5.1 RQ1: Tuning method performance
Methods2Testrunnable HumanEval-Xjava
@@
Instr. Cov.
Branch Cov.
Mutation Score
Valid syntax
CodeBLEU
pass@1
[…]
(a) (b) (c)
Method Trainable
params
Methods2Testrunnable HumanEval-Xjava
Valid syntax
CodeBLEU
pass@1
Instr. Cov.
Branch Cov.
Mutation Score
Valid syntax
CodeBLEU
pass@1
[…]
None 0 0.96 0.17 0.09 0.52 0.12 0.23 1.0 0.28 0.03
Fine-tuning 356,712,448 0.98 0.22 0.14 (0.48) 0.16 0.83 1.0 (0.24) (0.02)
LoRA 1,310,720 0.96 0.19 0.15 (0.48) 0.15 0.41 1.0 0.29 (0.02)
(IA)3 143,360 (0.95) 0.17 0.13 (0.5) 0.13 0.45 1.0 0.28 0.03
Prompt tuning 20,480 0.96 0.17 0.12 (0.43) 0.13 0.44 1.0 (0.25) (0.02)
CodeGen2-1B
None 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Fine-tuning 1,015,306,240 0.75 0.1 0.04 0.31 0.09 0.11 0.05 0.01 0.0
LoRA 2,097,152 0.38 0.03 0.12 0.01 0.0 0.74 0.09 0.01 0.0
(IA)3 229,376 0.02 0.26 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Prompt tuning 40,960 0.7 0.26 0.0 0.0 0.0 0.0 0.08 0.25 0.0
StarCoder2-3B
None 0 0.93 0.1 0.53 0.14 0.06 0.52 1.0 0.41 0.22
Fine-tuning 3,030,371,328 0.96 0.23 (0.17) 0.62 0.22 (0.38) 1.0 0.41 (0.19)
LoRA 4,546,560 0.98 0.24 (0.19) 0.59 0.2 0.53 1.0 (0.32) (0.16)
(IA)3 468,480 0.94 0.21 (0.27) 0.41 0.14 (0.39) 1.0 0.42 0.24
Prompt tuning 61,440 0.93 0.1 0.54 0.15 0.06 (0.28) 1.0 (0.38) (0.18)
Qwen2.5-Coder-3B
None 0 0.97 0.21 0.18 0.53 0.19 0.0 1.0 0.42 0.34
Fine-tuning 3,085,938,688 0.98 0.25 (0.15) 0.53 (0.13) 0.0 1.0 (0.39) (0.21)
LoRA 3,686,400 0.98 0.23 0.2 0.55 0.19 0.51 1.0 0.45 0.4
(IA)3 479,232 0.97 0.22 0.19 0.6 0.22 0.64 1.0 0.46 0.38
Prompt tuning 40,960 0.97 0.22 0.2 0.53 0.2 0.55 (0.99) (0.36) (0.31)
CodeGen2-3.7B
None 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Fine-tuning 3,641,174,016 0.41 0.07 0.05 0.33 0.15 0.0 0.74 0.26 0.0
LoRA 4,194,304 0.42 0.07 0.12 0.09 0.06 0.58 0.4 0.04 0.0
(IA)3 458,752 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Prompt tuning 81,920 0.22 0.26 0.0 0.0 0.0 0.0 0.0 0.0 0.0
CodeLlama-7B
None 0 0.98 0.24 0.18 0.6 0.23 0.14 0.99 0.4 0.21
Fine-tuning 6,738,546,688 0.98 0.24 0.22 0.62 (0.22) (0.0) 1.0 0.41 0.22
LoRA 8,388,608 0.98 0.26 0.22 0.61 0.24 0.51 0.99 (0.35) (0.2)
(IA)3 614,400 0.98 0.24 0.2 (0.59) 0.24 0.36 1.0 (0.39) 0.21
Prompt tuning 81,920 (0.97) (0.23) (0.11) (0.59) (0.13) 0.19 0.99 (0.37) (0.2)
CodeGen2-7B
None 0 0.99 0.22 0.18 0.57 0.21 0.35 1.0 0.38 0.13
Fine-tuning 6,862,858,240 (0.98) 0.23 (0.15) (0.51) (0.2) (0.27) 1.0 (0.35) (0.08)
LoRA 8,388,608 0.99 0.24 (0.17) 0.6 0.21 0.41 1.0 (0.37) (0.12)
(IA)3 917,504 (0.98) 0.22 0.18 0.58 0.22 0.52 1.0 0.39 0.13
Prompt tuning 81,920 (0.98) 0.22 0.19 (0.54) 0.21 0.51 1.0 (0.36) (0.12)
StarCoder2-7B
None 0 0.92 0.11 0.53 0.12 0.05 0.37 1.0 0.35 0.23
Fine-tuning 7,173,923,840 0.97 0.24 (0.2) 0.6 0.2 (0.14) 1.0 0.43 (0.22)
LoRA 7,340,032 0.97 0.24 (0.21) 0.56 0.2 0.53 1.0 0.42 (0.22)
(IA)3 753,664 0.95 0.24 (0.2) 0.62 0.22 0.41 1.0 0.42 (0.21)
Prompt tuning 92,160 0.93 0.25 (0.2) 0.54 0.19 0.43 1.0 0.42 0.24
Qwen2.5-Coder-7B
None 0 0.97 0.23 0.2 0.59 0.21 0.49 1.0 0.39 0.37
Fine-tuning 7,615,616,512 (0.94) (0.19) (0.14) (0.57) (0.18) (0.0) 1.0 0.41 (0.18)
LoRA 5,046,272 0.98 0.25 0.23 0.62 0.21 0.51 1.0 (0.36) (0.32)
(IA)3 645,120 0.98 0.24 0.24 0.64 0.21 (0.39) 1.0 0.39 0.38
Prompt tuning 71,680 (0.96) 0.23 0.21 0.63 0.22 (0.36) 1.0 0.41 0.41
Qwen2.5-Coder-14B
None 0 0.98 0.25 0.24 0.58 0.2 0.4 1.0 0.37 0.3
Fine-tuning 14,770,033,664 (0.96) (0.24) (0.2) (0.55) (0.17) (0.12) 1.0 0.43 0.35
LoRA 12,502,912 (0.97) 0.26 0.24 0.61 0.21 0.53 1.0 0.41 0.38
(IA)3 958,464 0.98 0.26 0.24 0.61 0.2 (0.27) 1.0 0.39 0.34
Prompt tuning 102,400 0.98 0.25 (0.2) 0.59 0.2 0.54 1.0 0.39 0.38
StarCoderBase
None 0 0.92 0.11 0.48 0.15 0.06 0.28 0.99 0.4 0.2
Fine-tuning 15,517,456,384 0.97 0.26 (0.22) 0.58 0.22 0.41 1.0 (0.39) 0.22
LoRA 8,028,160 0.92 0.11 0.48 0.15 0.06 0.47 0.99 (0.37) 0.2
(IA)3 1,239,040 0.92 0.11 0.48 0.16 0.06 0.46 0.99 0.4 0.2
Prompt tuning 122,880 (0.91) 0.13 (0.34) 0.2 0.09 (0.25) (0.78) (0.22) (0.1)
StarCoder2-15B
None 0 0.93 0.13 0.53 0.17 0.07 0.37 1.0 0.34 0.25
Fine-tuning 15,655,899,136 0.98 0.25 (0.19) 0.49 0.17 (0.03) (0.99) 0.4 (0.23)
LoRA 12,124,160 0.98 0.26 (0.21) 0.61 0.22 (0.16) 1.0 0.38 0.26
(IA)3 1,249,280 0.93 0.13 0.53 0.17 0.07 0.43 1.0 0.34 0.25
Prompt tuning 122,880 0.98 0.26 (0.18) 0.62 0.24 0.52 1.0 0.44 0.28
CodeGen2-16B
None 0 0.98 0.23 0.18 0.57 0.22 0.7 1.0 0.39 0.13
Fine-tuning 16,032,155,648 0.99 0.24 (0.13) (0.51) (0.15) (0.05) (0.98) (0.28) (0.07)
LoRA 13,369,344 0.99 0.25 (0.17) (0.52) (0.19) (0.29) 1.0 (0.37) 0.14
(IA)3 1,462,272 0.98 0.23 0.19 0.57 (0.21) (0.53) 1.0 0.39 0.13
Prompt tuning 122,880 0.98 0.23 0.19 0.61 0.24 (0.54) 1.0 (0.37) 0.13
5.1.1 Syntactic Correctness
@@
As shown in Section 5.1(b), the percentage of syntactically correctly generated unit tests is relatively high across the models, with >> 91% for all except the smaller CodeGen2-1B and CodeGen2-3.7B models (marked in red in Section 5.1). These two cannot produce any valid unit tests with the base model. However, after tuning, both models are able to produce up to 75% and 42% syntactically valid tests, respectively. Notable findings from Section 5.1(b)) are as follows: 1) The increase of 70% using prompt tuning of CodeGen2-1B, tuning only 0.02‰ (40 thousand out of 1 billion parameters) of the parameters. 2) The increase of 42% using LoRA on CodeGen2-3.7B, tuning only 0.1% (4.19 million out of 3.7 billion parameters) of the parameters.
@@
Further, data in Section 5.1(b) illustrate that: 1) Full fine-tuning achieves higher percentages of syntactically correct unit tests than the baseline. The median value of syntactically valid tests for the full fine-tuning is 0.97, which is 1% higher than the value of the baseline (0.96). 2) There is a slight tendency for models with more trainable parameters to generate a higher proportion of syntactically correct unit tests. 3) The difference in syntactical validity between fine-tuning and PEFTs is not very prominent. The median values of syntactically valid tests of PEFT methods are: LoRA 0.98, (IA)3 0.95, and prompt tuning 0.96. However, considering that full fine-tuning often tunes parameters several thousand more times than PEFT methods, it is probably not wise to use full fine-tuning to improve the syntactic correctness of the unit tests if cost is a concern.
5.1.2 Similarity
The Methods2Testrunnable dataset CodeBLEU column in Section 5.1(b) shows the CodeBLEU scores of the baseline, full fine-tuning, and PEFT methods. Due to the low number of syntactically valid tests generated from the CodeGen2-1B and CodeGen2-3.7B models, we exclude these models from further analysis. The median CodeBLEU scores are: baseline 0.21, fine-tuning 0.24, LoRA 0.24, (IA)3 0.22, and prompt tuning 0.23. The notable results are: 1) LoRA achieve the highest CodeBLEU scores in seven of the eleven valid models, improving the median CodeBLEU value by 14% over the baseline. 2) Prompt tuning is also the only PEFT method that shows slight negative results.
5.1.3 Passing rate
Table 5: Test execution statuses for StarCode2-7B.
Method Succ. Failed Interrupt CompErr NoAssert
None 53% 3% 3% 16% 26%
Fine-tuning 20% 14% 9% 51% 6%
LoRA 21% 15% 9% 51% 5%
(IA)3 20% 14% 10% 49% 7%
Prompt tuning 20% 15% 7% 52% 6%
As explained in section 4.8, we consider a test as passing if it executes successfully without errors, as determined by a test oracle. Any tests that did not run due to an environmental exception (see section 4.8) are excluded from analysis because such cases did not reflect the quality of the generated test cases. As we use Methods2Testrunnable to generate and evaluate tests, the original focal method under test serves as a proxy for the oracle, effectively acting as the ground truth. While we do not assume that the methods under test are bug-free, we trust that all methods from the Methods2Test dataset are at least passable and thus suitable as oracles for evaluating generated tests. However, using the method as the oracle of test passing introduces a limitation: tests performing no assertion checks may still return passing results, which are misleading. To address this issue, we heuristically mark the generated tests as failed (even if their executions return passing results) if they do not contain any of the substrings “assert”, “verify”, or “fail”, which are indicative of meaningful assertions or validation logic in the test body.
From the pass@1 column in Section 5.1(b), we calculated individual median pass@1 values and got the results: baseline 0.20, fine-tuning 0.17, LoRA 0.21, (IA)3 0.20, and prompt tuning 0.20. The baseline achieves the highest pass@1 rate on five of the eight models, with a significantly higher median score than all tuning methods. To understand this phenomenon, we analyzed and categorized different reasons for tests not returning a passing result. Table 5 shows an example of statistical analysis results of the execution status of test cases generated by the StarCode2-7B model. We found that the primary cause of the low number of successful test executions after fine-tuning is compilation errors. Although most of the generated content is syntactically valid (see section 5.1.1), compilation errors still happen, mainly due to the generated content calling nonexistent methods or having type mismatches. Most of these compilation errors stem from repetition or boilerplate patterns that superficially satisfy the heuristic of unit test generation without actually validating the method’s behavior.
5.1.4 Coverage
@@
Among the tests returning passing results, the individual median values of the instruction coverage are: baseline 0.53, fine-tuning 0.55, LoRA 0.59, (IA)3 0.58, and prompt tuning 0.54. The individual median values of the branch coverage are: baseline 0.19, fine-tuning 0.18, LoRA 0.20, (IA)3 0.21, and prompt tuning 0.20. These numbers show that the code coverage is generally high among the tuned models. It means that if the tuned models can create runnable tests, the tests are likely to cover a significant portion of the code to potentially identify edge cases. The results also show that LoRA and full fine-tuning outperform other methods, although the differences between their performance and other PEFT methods are insignificant.
5.1.5 Mutation score
Of the tests that pass and returning passing results, the individual median values of the mutation scores are: baseline 0.37, fine-tuning 0.12, LoRA 0.51, (IA)3 0.43, and prompt tuning 0.44. A notable finding is that full fine-tuning underperforms the baseline, suggesting that extensive parameter updates may diminish the model’s ability to generate fault-revealing tests. Conversely, all PEFT methods outperform the baseline. LoRA is the most performant PEFT method in five of the eleven models, improving the median mutation scores value by 38% over the baseline.
5.1.6 Resilience against catastrophic forgetting
@@
The results in Section 5.1(c) show the performance of the baseline and tuned models on a normal code generation task using the Java subset of the HumanEval-X dataset. The results show some cases of degradation (marked using parentheses in Section 5.1(c)) of code generation when comparing the results of the tuned models with the baseline model. This applies both to full fine-tuning and PEFT. However, the levels of degradation are, for the most part, trivial and inconsistent. The median CodeBLEU scores are: baseline 0.39, fine-tuning 0.40, LoRA: 0.37, (IA)3 0.39, and prompt tuning 0.37. The median pass@1 values are: baseline 0.22, fine-tuning 0.21, LoRA 0.20, (IA)3 0.21, and prompt tuning 0.20. In several cases, e.g. StarCoder2-15B, PEFT tuning actually improves the CodeBLEU scores and pass@1 value. This shows that tuning methods are mostly resilient against catastrophic forgetting, and PEFT methods exhibit transfer learning capabilities comparable to those achieved through full fine-tuning.
Summary for RQ1: PEFT could be a better choice than full fine-tuning if cost-effectiveness to improve the syntactic correctness of the generated unit tests is a concern. LoRA can generate test cases as good as full fine-tuning, with consistently better mutation score. The tuned models may generate fewer executable test cases than baseline models, although the executable ones are better at test coverage. Further, PEFT methods are mostly resilient against catastrophic forgetting.
5.2 RQ2: PEFT training resource utilization effectiveness